OpenStudy (anonymous):

Find the slope of the line passing through the points (4,-2) and (-1,1) and put the equation in standard form. (Ax + By = C) Please show all your work so i can do the next one myself.

OpenStudy (ranga):

Slope = (y2-y1) / (x2-x1)

OpenStudy (ranga):

Once you know the slope m, (y-y1) = m(x-x1) is the equation of the line. You can rearrange the terms to put it in standard form.

OpenStudy (ranga):

The points are (4,-2) and (-1,1) Slope m = (y2-y1) / (x2-x1) = (1 - (-2)) / (-1 - 4) = (1+2)/(-5) = -3/5 Knowing the slope m and any point on the line (either (4,-2) or (-1,1)) we can find the equation of the line. We will choose the (-1,1) point because it is simpler than (4,-2): y - y1 = m(x - x1) y - 1 = -3/5(x - (-1)) = -3/5(x + 1) y - 1 = -3/5x - 3/5 y = -3/5x - 3/5 + 1 y = -3/5x + 2/5 multiply throughout by 5 5y = -3x + 2 add 3x to both sides: 3x + 5y = 2 is the equation of the line in standard form.

OpenStudy (ranga):

After find the equation, it is a good idea to put the points back and see if they satisfy the equation. I tried both (4, -2) and (-1, 1) in 3x + 5y = 2 and both satisfy the equations.